OpenStudy (anonymous):

Need help finishing this code its supposed to print out commas after every third # but it is adding an extra 0 at the end of both commas

OpenStudy (anonymous):

#include<iostream> using namespace std; int main() { int first3, middle3, last3; int original, temp, first, second,third, fourth, first1, second1, third1, fourth1; cout<<"enter a 7-9 digit number"<<endl; cin>>original; temp=original/1000; last3=original%1000; first=last3/10; second=last3%10; third=first/10; fourth=first%10; first3=temp/1000; middle3=temp%1000; first1=middle3/10; second1=middle3%10; third1=first1/10; fourth1=first1%10; cout<<"the number with a comma is "<<endl; cout<<first3<<","<<first1<<second1<<third1<<","<<first<<second<<third<<endl; return 0; }

OpenStudy (anonymous):

i doesnt put zeros it just puts some random numbers

OpenStudy (espex):

http://ideone.com/RhtIN input: 123456789 output: enter a 7-9 digit number temp= 123456 last3= 789 first= 78 second= 9 third= 7 fourth= 8 first3= 123 middle3= 456 first1= 45 second1= 6 third1= 4 fourth1= 5 the number with a comma is 123,4564,7897 It is just a matter of keeping track of your variables.

OpenStudy (anonymous):

@eSpeX So what would i type in the last cout statement to get it write

OpenStudy (espex):

cout<<first3<< ","<<middle3<<","<< last3<< endl; I added this line to your code and below is the output with your code and the new code as the last two lines. input: 1234567 output: enter a 7-9 digit number temp= 1234 last3= 567 first= 56 second= 7 third= 5 fourth= 6 first3= 1 middle3= 234 first1= 23 second1= 4 third1= 2 fourth1= 3 the number with a comma is 1,2342,5675 1,234,567 http://ideone.com/8KhZZ

OpenStudy (rsmith6559):

A pretty printed number is a string. it's easier to do the fomatting with the string than th number.